Total combinations for drawing 5 cards is 52 choose 5 which is...52*51*50*49*48/(5*4*3*2*1)
possible outcomes for drawing 5 hearts is 13 choose 5 which is...13*12*11*10*9/(5*4*3*2*1)
Probability of drawing 5 hearts is (52 choose 5)/(13 choose 5) which is....000495
How to make a working model on angles of parallel lines?
huh... i also askd dis question... bt no answer was der...:(
When rolling a pair of dice is there a better chance of rolling a seven or a nine?
There is a better chance of rolling a 7 because there are more combinations of numbers on a pair of dice that add to 7 (the probability of rolling a 7 is 6/36 vs a 9 which is 4/36).
1 2 3 4 5 6
+-------------------
1 | 2 3 4 5 6 7
2 | 3 4 5 6 7 8
3 | 4 5 6 7 8 9
4 | 5 6 7 8 9 10
5 | 6 7 8 9 10 11
6 | 7 8 9 10 11 12
What is the probability of rolling two dice and it being composite?
There are 36 possible outcomes when two dice are rolled.
The composite outcomes and the number of corresponding events are:
4 - 3
6 - 5
8 - 5
9 - 4
10 - 3
12 - 1
So there are 21 outcomes out of 36 where the result is a composite.
So probability of a composite = 21/36 = 7/12 = 0.58333 or 58.333... %
What is the probability of rolling a 3 and 6 on a six sided dice?
The probability of rolling a 3 is 1/6. The probability of rolling a 6 is 1/6.
There are two ways to roll a 3 and 6: roll 3 on the first die and 6 on the second or roll 6 on the first and 3 on the second. The probability of the first is 1/6 x 1/6 = 1/36. Likewise the probability of the second is 1/6 x 1/6 = 1/36. Combined you get 1/36 + 1/36 = 2/36 = 1/18
What is the probability of picking a non face card?
Approximately 77% that you will not choose a face card
8 ways
Is it true that two dependent events can have the same probability of occurring?
Yes, it is possible for two dependent events to have the same probability of occurring. The probability of an event is dependent on the outcomes of other events, and it is influenced by the relationship between these events. So, it is conceivable for two dependent events to have equal probabilities.
Does 60 percent of rain means its going to rain?
No. A 60% chance of rain means exactly what it sounds like: there is a 60% chance that it will rain. This also means that there is a 40% chance that it will not rain.
What does The Right of Passage mean?
answer: A ritual or ceremony signifying an event in a person's life indicative of a transition from one stage to another, as from adolescence to adulthood or even an initiation?
State and discuss limitations of statistics?
Despite The usefulness of statistics in Many fields, impression should not be carried that statistics are like magical devices which always provide the correct solution of problems.
1. Statistics does not deal with isolated measurement:not all quantitative data are statistical. Isolated measurements are not also statistical. Data are statistical when they are related to measurement of masses, not statistical when they are related to an individual item or event as a separate entity.
2. Statistics deals only with quantitative characteristics: statistics are numerical statements of facts. Such characteristics cannot expressed in numbers are incapable of statistics analysis.
3. statistical results are true only on an average: the conclusion obtained statistically are not universally true; they are true only under certain conditions.
4. Statistics is only a means: Statistical methods furnish only one method of studying a problem. They may not provide the vest solution under all circumstances.
5. Statistics can be misused: The greatest limitation of statistics is that it is liable to be misused. The misuse of statistics may arise because of several reasons. For example, if statistical conclusion are based on incomplete information, one may arrive at fallacious conclusions.
What is the jack in a deck of playing cards?
Well it depends what your playing it could be a 5 or a 10. It could be a lot of numbers. In Poker its probually 11.
Why is probability an exact science?
The origin of logic comes from scientific deduction of thought. The only science that is exact, is the science and facts that cannot be known for sure.or at least seen with a physical eye. In the beginning God created the heavens and the earth. So the probability of an exact science is... inconceivable.
If two dice were rolled what is the probability of a sum of 5 being rolled?
1 in 9. Since there are 36 (6x6) possible combinations, and 4 of them (1+4, 4+1, 2+3, 3+2) will produce a sum of 5, the probability is 4 in 36, or 1 in 9.
What is probability mass function?
It is a function which is usually used with continuous distributions, to give the probability associated with different values of the variable.
Alpha is the significance level of the test and has little to do with the failure rate of parts (or even parst!).
What is the probability it will snow this Christmas?
There is insufficient information in the question to properly answer it. The probability of snow at any given time depends on location, which you did not give. Please restate the question.
What is the probability of tossing tails and rolling a 2?
Tossing tails: 1 in 2 or 1/2
Rolling a 2: 1 in 6 or 1/6